OpenStudy (anonymous):

square root 90 + square root 10

OpenStudy (anonymous):

Is it your question ?: \[\sqrt{90}+\sqrt{10}\] K ??? Can you write it ?:)

OpenStudy (anonymous):

@lcasazza3

OpenStudy (anonymous):

yes, it is..I get 4sqrt10, not sure I am correct

OpenStudy (anonymous):

\[\sqrt{90}+\sqrt{10}=(\sqrt{9}*\sqrt{10})+\sqrt{10}\]\[\sqrt{10}(\sqrt{9}+1)=\sqrt{10}(3+1)=4\sqrt{10}\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

No problem!